The sum of two numbers is 6 less than twice the first number. Their difference is 10 less than four times the second number. Find each of the numbers.
Answer by Maths68(1474)

Find each of the numbers.
Let
First Number = x
Second Number = y
Given
The sum of two numbers is 6 less than twice the first number.
x+y=2x-6
x-2x=-y-6
-x=-y-6
Multiply by -1 both sides
(-1)(-x)=(-1)(-y-6)
x=y+6
x-y=6..................(1)

Their difference is 10 less than four times the second number.
x-y=4y-10
x-y-4y=-10
x-5y=-10...............(2)
Subtract (2) from (1)
x-5y=-10
-x+y=-6
-------------
-4y=-16
-4y/-4=-16/-4
y=4
Put the value of y in (1)
x-y=6..................(1)
x-4=6
x=6+4
x=10


First Number = x = 10
Second Number = y = 4